# Rare and exclusive few-body decays of the Higgs, Z, W bosons, and the top quark

### v3

@@ III.3 Radiative Higgs leptonium decays
All branching fractions computed here for Higgs decays into Z or $\gamma$ plus leptonium are listed in Table 8 and shown in graphical form in Fig. 10. As aforementioned, the decay rates are minuscule, in the $\mathcal{O}(10^{-11}$$10^{-16})$ range, and only BSM effects enhancing them can made them visible. If higher-order corrections increase the $\rm H\to\gamma+(\mathrm{e}^{+}\mathrm{e}^{-})_{1}$ channel by about 400 times as indicated in Ref. [121], then a machine producing as many Higgs bosons as the FCC-hh can attempt to observe the $\rm H\to\gamma+(3\gamma)$ with a displaced triphoton vertex from the late orthopositronium $(ee)_{1}\to 3\gamma$ decay, or $\rm H\to\gamma+\mathrm{e}^{+}\mathrm{e}^{-}$ with a displaced dielectron from the magnetic-field-induced breaking of orthopositronium into its constituents.
Table 8: Compilation of exclusive Higgs decays to a photon or a Z boson plus an ortho- $(\ell^{+}\ell^{-})_{1}$ or para- $(\ell^{+}\ell^{-})_{0}$ leptonium state (only the ground states, $n=1$, are considered). For each decay, we provide the prediction of its branching fraction computed here. The last two columns indicate whether the decay can be produced at FCC-ee/FCC-hh.
@@
$\rm H\to\,V$ $+$ $\rm(\ell\ell)$ Branching fraction Framework 2024 HL-LHC FCC-ee FCC-hh
$\rm H\,\to\,\gamma$ + $\rm(ee)_{1}$  $3.5$$\times~{}10^{-12}$ (this work)
$\rm(\mu\mu)_{1}$  $3.5$$\times~{}10^{-12}$ (this work)
$\rm(\tau\tau)_{1}$  $2.2$$\times~{}10^{-12}$ (this work)
$\rm H\,\to\,Z$ + $\rm(ee)_{1}$  $5.2$$\times~{}10^{-13}$ (this work)
$\rm(\mu\mu)_{1}$  $5.7$$\times~{}10^{-13}$ (this work)
$\rm(\tau\tau)_{1}$  $1.4$$\times~{}10^{-11}$ (this work)
$\rm(ee)_{0}$  $2.7$$\times~{}10^{-16}$ (this work)
$\rm(\mu\mu)_{0}$  $1.1$$\times~{}10^{-14}$ (this work)
$\rm(\tau\tau)_{0}$  $3.2$$\times~{}10^{-12}$ (this work)
@@ III.4 Exclusive Higgs decays into two mesons
Figure 11 shows representative diagrams of the exclusive decay of the Higgs boson into two mesons, which can proceed through a multitude of intermediate states coupling to the scalar particle: quarks, gluons, virtual EW gauge bosons. First estimates of these processes were performed ignoring the internal motion of the produced quark-antiquark pairs [123], and then further improved within different approaches for the meson-pair formation [124, 125, 126, 127, 128, 129]. Except for a channel involving the $\phi$ meson, only exclusive decays involving charmonium and/or bottomonium final states have been computed to date, i.e., no calculations of decays to a pair of light mesons exist to our knowledge (due to the difficulties explained in Section IV.4 for the similar $\rm Z\to M+M$ case).
@@
Table 9 lists the corresponding theoretical predictions and experimental limits for concrete $\rm H\to 2(\mathrm{Q}\overline{\mathrm{Q}})$ and $\rm H\to(\mathrm{Q}\overline{\mathrm{Q}})(\mathrm{Q}\overline{\mathrm{Q}}^{\prime})$ decay modes. The calculations are carried out in multiple frameworks (LC + LCDA, RQM, NRQCD/NRCSM, NRQCD + LDME) and predict rates in the $\mathcal{O}(10^{-9}$$10^{-11})$ range [^6] Semiexclusive decays $\rm H\to B_{\mathrm{c}}+Q,+Q^{\prime},\,B_{\mathrm{c}}^{*}+Q+Q^{\prime}$ have much larger rates [130] and, since they are not rare nor exclusive, they are not covered here., with some differences in the results for the same final state driven by the partial inclusion of the diagrams shown in Fig. 11. As a matter of fact, the existing predictions have not fully included all processes shown in the figure with, e.g., the W-induced and quark “crossed” decays often considered subleading and not added to the rates. The $\rm H\to\phi+\mathrm{J}/\psi$ decay is the only process computed to date that includes the $\rm H\to W^{*}W^{*}$ intermediate diagram. The direct (quark-induced) contributions are only relevant for the heavier double bottomonia with larger Yukawa couplings, and calculations of final decays involving charmonium ignore them.
Table 9: Compilation of exclusive Higgs decays to two mesons. For each decay, we provide the predicted branching fraction(s) and the theoretical approach used to compute it, as well as the current experimental upper limit and that estimated for HL-LHC. The last two columns indicate whether the decay can be produced at FCC-ee/FCC-hh.
@@
$\rm H\,\to$ M + M Branching fraction Framework 2024 HL-LHC FCC-ee FCC-hh
$\rm H\,\to$ $\rm\phi$ + $\rm J/\psi$  $1.0$$\times~{}10^{-9}$ LC+LCDA [124]
$\rm J/\psi$ + $\rm J/\psi$  $1.7$$\times~{}10^{-10}$ RQM [126] $<$ $3.8$×$\times~{}10^{-4}$ [107] $\lesssim 5.8\times~{}10^{-5}$
@@
 $2.1$$\times~{}10^{-10}$ RQM [128]
 $(5.9\pm 2.3)$$\times~{}10^{-10}$ NRQCD/NRCSM [127]
 $1.5$$\times~{}10^{-10}$ LC+LCDA [124]
$\rm\psi(2S)$ + $\rm J/\psi$  $\mathcal{O}(5)$$\times~{}10^{-11}$ $<$ $2.1$×$\times~{}10^{-3}$ [107] $\lesssim 3.2\times~{}10^{-4}$
$\rm\psi(2S)$  $(5.1\pm 2.0)$$\times~{}10^{-11}$ NRQCD/NRCSM [127] $<$ $3.0$×$\times~{}10^{-3}$ [107] $\lesssim 4.5\times~{}10^{-4}$
@@
 $1.6$$\times~{}10^{-11}$ LC+LCDA [124]
$\rm\Upsilon(1S)$  $1.8$$\times~{}10^{-10}$ RQM [126] $<$ $1.7$×$\times~{}10^{-3}$ [107] $\lesssim 2.6\times~{}10^{-4}$
[…]
 $2.3$$\times~{}10^{-9}$ RQM [128]
 $(4.3\pm 0.9)$$\times~{}10^{-10}$ NRQCD/NRCSM [127]
 $2.3$$\times~{}10^{-9}$ LC+LCDA [124]
$\rm\Upsilon(2S)$ + $\rm\Upsilon(2S)$  $(1.0\pm 0.2)$$\times~{}10^{-10}$ NRQCD/NRCSM [127]
$\rm\Upsilon(3S)$ + $\rm\Upsilon(3S)$  $(5.7\pm 1.2)$$\times~{}10^{-11}$ NRQCD/NRCSM [127]
$\rm\Upsilon(mS)$ + $\rm\Upsilon(nS)$ $<$ $3.5$×$\times~{}10^{-4}$ [107] $\lesssim 9.2\times~{}10^{-6}$ [61]
Experimentally, the double-$(\mathrm{Q}\overline{\mathrm{Q}})$ and $(\mathrm{Q}\overline{\mathrm{Q}})(\mathrm{Q}\overline{\mathrm{Q}}^{\prime})$ decays have been searched for at the LHC [107], but the current $\mathcal{O}(10^{-3}$$10^{-4})$ limits are 4–5 orders-of-magnitude larger than the predictions. Their production at the HL-LHC, as well as at any future lepton collider, appears unfeasible, and only a machine like FCC-hh will have the Higgs production rates required to reach those decay modes, as indicated in Fig. 12.
@@ IV.1 Rare three-body Z boson decays
The two-body decay of the Z boson into two massless vector particles ($\rm Z\to\gamma\gamma,gg$) is forbidden by the Landau–Yang theorem, which states that a spin-1 boson cannot decay into two massless bosons [131, 132], and of course the $\rm Z\to\gamma g$ decay is forbidden by conservation of colour. On the other hand, three-body decays into gauge bosons, or into a photon plus two neutrinos, are possible but can only be induced at the loop level in the SM (Fig. 13). Calculations for these virtual processes exist for many years [133, 134, 135, 136] but they are not all always consistent among each other and, to our knowledge, have not been revised since then, although key underlying parameters, such as the $\alpha_{\text{s}}(m_{{}_{\mathrm{Z}}})$ coupling constant, hadronic Z boson partial widths, and top quark mass, have (noticeably) changed. For example, the original $\rm Z\to 3\gamma$ calculations from [134] used $m_{\mathrm{t}}=124$ GeV, whereas the calculation of the $\rm Z\to 3g$ decay of [135] used the $m_{\mathrm{b}}\to 0$, $m_{\mathrm{t}}\to\infty$ limits, and the prediction for $\rm Z\to g+g+\gamma$ from [133] used $\alpha_{\text{s}}(m_{{}_{\mathrm{Z}}})=0.17$, $m_{\mathrm{t}}=20$ GeV, and $\Gamma_{\mathrm{Z}}=\Gamma_{\mathrm{Z\to\mathrm{q}\overline{\mathrm{q}}}}\approx 2$ GeV. We note also that many of these older works, including the review of Ref. [18], provide contradictory values for some of these different branching fractions. We have therefore recalculated the rates for all these Z boson rare decays with MG5_aMC in the SM framework with QCD and EW loop corrections computed up to NLO accuracy, with massive fermions in the loops where needed, using the input parameters of Table 2. The updated branching fractions (Table 10) change with respect to the existing ones (using the old formulas with updated SM parameters, where needed) by 10–25% up or down, except for the $\rm Z\to\gamma+\nu\overline{\nu}$ case, where our result is six times smaller than that estimated in [136]. This older work quoted a partial width for the triangle diagram which is smaller than our results for this contribution, whereas the relative impact of box diagrams was not clearly specified. So we suspect that a larger cancellation of contributions is present in our setup compared to that in this previous study.
@@
Branching fraction Framework 2024 HL-LHC FCC-ee
$\rm Z\;\to\;g+g+g$  $1.9$$\times~{}10^{-6}$ NLO SM [135] $<$ $1.1$×$\times~{}10^{-2}$ [137]
 $1.8$$\times~{}10^{-6}$ NLO SM, MG5_aMC (this work)
$\rm Z\;\to\;g+g+\gamma$  $7.0$$\times~{}10^{-7}$ NLO SM [^7] With updated SM parameters. The original result was $\mathcal{B}=1.8\times 10^{-6}$ [133] [133]
 $6.9$$\times~{}10^{-7}$ NLO SM, MG5_aMC (this work)
$\rm Z\;\to\;g+g+\gamma_{\mathrm{soft}}$  $1.2$$\times~{}10^{-9}$ NLO SM, MG5_aMC (this work)
$\rm Z\;\to\;\gamma+\gamma+\gamma$  $5.4$$\times~{}10^{-10}$ NLO SM [134] $<$ $2.2$×$\times~{}10^{-6}$ [66] $\lesssim 1.3\times~{}10^{-7}$
 $6.6$$\times~{}10^{-10}$ NLO SM, MG5_aMC (this work)
$\rm Z\;\to\;\gamma+\gamma+\gamma_{\mathrm{soft}}$  $2.2$$\times~{}10^{-12}$ NLO SM, MG5_aMC (this work) $<$ $1.5$×$\times~{}10^{-5}$ [138] $\lesssim 2.1\times~{}10^{-7}$
$\rm Z\;\to\;\gamma+\nu+\hat{\nu}$  $7.2$$\times~{}10^{-10}$ NLO SM [136]
 $1.4$$\times~{}10^{-10}$ NLO SM, MG5_aMC (this work)
@@
Figure 14: Branching fractions (in negative log scale) of rare three-body Z boson decays. The most recent theoretical predictions (blue bars) are compared to current experimental upper limits (violet) and expected conservative HL-LHC bounds (orange). The red vertical line indicates the expected FCC-ee reach based only on the total number of Z bosons to be produced.

### v4

@@ III.3 Radiative Higgs leptonium decays
All branching fractions computed here for Higgs decays into Z or $\gamma$ plus leptonium are listed in Table 8 and shown in graphical form in Fig. 10. As aforementioned, the decay rates are minuscule, in the $\mathcal{O}(10^{-11}$$10^{-16})$ range, and only BSM effects enhancing them can made them visible. If higher-order corrections increase the $\rm H\to\gamma+(\mathrm{e}^{+}\mathrm{e}^{-})_{1}$ channel by about 400 times as indicated in Ref. [Martynenko:2024rfj], then a machine producing as many Higgs bosons as the FCC-hh can attempt to observe the $\rm H\to\gamma+(3\gamma)$ with a displaced triphoton vertex from the late orthopositronium $(ee)_{1}\to 3\gamma$ decay, or $\rm H\to\gamma+\mathrm{e}^{+}\mathrm{e}^{-}$ with a displaced dielectron from the magnetic-field-induced breaking of orthopositronium into its constituents.
Table 8: Compilation of exclusive Higgs decays to a photon or a Z boson plus an ortho- $(\ell^{+}\ell^{-})_{1}$ or para- $(\ell^{+}\ell^{-})_{0}$ leptonium state (only the ground states, $n=1$, are considered). For each decay, we provide the prediction of its branching fraction computed here. The last two columns indicate whether the decay can be produced at FCC-ee/FCC-$h$.
@@
[…]
$\rm H\,\to\,\gamma$ + $\rm(ee)_{1}$ $3.5$ $\times~10^{-12}$ Eq. (25) (this work)
$1.1$ $\times~10^{-11}$ RQM [Martynenko:2024rfj]
$\rm(\mu\mu)_{1}$ $3.5$ $\times~10^{-12}$ Eq. (25) (this work)
$1.1$ $\times~10^{-11}$ RQM [Martynenko:2024rfj]
$\rm(\tau\tau)_{1}$ $2.2$ $\times~10^{-12}$ Eq. (25) (this work)
$3.5$ $\times~10^{-12}$ RQM [Martynenko:2024rfj]
$\rm H\,\to\,Z$ + $\rm(ee)_{1}$ $5.2$ $\times~10^{-13}$ Eq. (29) (this work)
$7.9$ $\times~10^{-13}$ RQM [Martynenko:2024rfj]
$\rm(\mu\mu)_{1}$ $5.7$ $\times~10^{-13}$ Eq. (29) (this work)
$9.8$ $\times~10^{-13}$ RQM [Martynenko:2024rfj]
$\rm(\tau\tau)_{1}$ $1.4$ $\times~10^{-11}$ Eq. (29) (this work)
$5.7$ $\times~10^{-11}$ RQM [Martynenko:2024rfj]
$\rm(ee)_{0}$ $2.7$ $\times~10^{-16}$ Eq. (29) (this work)
$\rm(\mu\mu)_{0}$ $1.1$ $\times~10^{-14}$ Eq. (29) (this work)
$\rm(\tau\tau)_{0}$ $3.2$ $\times~10^{-12}$ Eq. (29) (this work)
@@ III.4 Exclusive Higgs decays into two mesons
Figure 11 shows representative diagrams of the exclusive decay of the Higgs boson into two mesons, which can proceed through a multitude of intermediate states coupling to the scalar particle: quarks, gluons, virtual EW gauge bosons. First estimates of these processes were performed ignoring the internal motion of the produced quark-antiquark pairs [Doroshenko:1987nj], and then further improved within different approaches for the meson-pair formation [Kartvelishvili:2008tz, Belov:2021toy, Faustov:2022jfk, Gao:2022iam, Faustov:2023phi, Belov:2023iop]. Except for a channel involving the $\phi$ meson, only exclusive decays involving charmonium and/or bottomonium final states have been computed to date, i.e., no calculations of decays to a pair of light mesons exist to our knowledge (due to the difficulties explained in Section IV.4 for the similar $\rm Z\to M+M$ case).
@@
Table 9 lists the corresponding theoretical predictions and experimental limits for concrete $\rm H\to 2(\mathrm{Q}\overline{\mathrm{Q}})$ and $\rm H\to(\mathrm{Q}\overline{\mathrm{Q}})(\mathrm{Q}\overline{\mathrm{Q}}^{\prime})$ decay modes. The calculations are carried out in multiple frameworks (LC + LCDA, RQM, NRQCD/NRCSM, NRQCD + LDME) and predict rates in the $\mathcal{O}(10^{-9}$$10^{-11})$ range [^8] Semiexclusive decays $\rm H\to B_{\mathrm{c}}+Q,+Q^{\prime},\,B_{\mathrm{c}}^{*}+Q+Q^{\prime}$ have much larger rates [Jiang:2015pah] and, since they are not rare nor exclusive, they are not covered here., with some differences in the results for the same final state driven by the partial inclusion of the diagrams shown in Fig. 11. As a matter of fact, the existing predictions have not fully included all processes shown in the figure with, e.g., the W-induced and quark “crossed” decays often considered subleading and not added to the rates. The $\rm H\to\phi+\mathrm{J}/\psi$ decay is the only process computed to date that includes the $\rm H\to W^{*}W^{*}$ intermediate diagram. The direct (quark-induced) contributions are only relevant for the heavier double bottomonia with larger Yukawa couplings, and calculations of final decays involving charmonium ignore them.
Table 9: Compilation of exclusive Higgs decays to two mesons. For each decay, we provide the predicted branching fraction(s) and the theoretical approach used to compute it, as well as the current experimental upper limit and that estimated for HL-LHC. The last two columns indicate whether the decay can be produced at FCC-ee/FCC-$h$.
@@
$\rm H\,\to$ M + M Branching fraction Framework 2024 HL-LHC FCC-$e$ FCC-$h$
$\rm H\,\to$ $\rm\phi$ + $\rm J/\psi$ $1.0$ $\times~10^{-9}$ LC+LCDA [Kartvelishvili:2008tz]
$\rm J/\psi$ + $\rm\eta_{c}$ $1.9$ $\times~10^{-15}$ RQM [Martynenko:2025nsg]
@@
$(5.8\text{ -- }6.0)$ $\times~10^{-9}$ NRQCD+LDME [Belov:2023iop]
$2.1$ $\times~10^{-10}$ RQM [Faustov:2023phi]
$(5.9\pm 2.3)$ $\times~10^{-10}$ NRQCD/NRCSM [Gao:2022iam]
$1.5$ $\times~10^{-10}$ LC+LCDA [Kartvelishvili:2008tz]
@@
$\rm\psi(2S)$ $(5.1\pm 2.0)$ $\times~10^{-11}$ NRQCD/NRCSM [Gao:2022iam] << $03\times~10^{-3}$ [CMS:2022fsq] $\lesssim 4.5\times~10^{-4}$
$\rm\chi_{c1}$ + $\rm\eta_{c}$ $1.9$ $\times~10^{-15}$ RQM [Martynenko:2025nsg]
$\rm h_{c}$ + $\rm J/\psi$ $1.9$ $\times~10^{-15}$ RQM [Martynenko:2025nsg]
$\rm B_{c}^{\mp}$ + $\rm B_{c}^{\pm}$ $(2.0\text{ -- }3.0)$ $\times~10^{-10}$ RQM [Belov:2021toy]
$\rm B_{c}^{*\mp}$ + $\rm B_{c}^{*\pm}$ $(1.4\text{ -- }1.7)$ $\times~10^{-10}$ RQM [Belov:2021toy]
$\rm\Upsilon(1S)$ + $\rm J/\psi$ $(2.7\text{ -- }3.6)$ $\times~10^{-10}$ NRQCD+LDME [Belov:2023iop]
$1.6$ $\times~10^{-11}$ LC+LCDA [Kartvelishvili:2008tz]
@@
$(8.5\text{ -- }9.2)$ $\times~10^{-10}$ NRQCD+LDME [Belov:2023iop]
$2.3$ $\times~10^{-9}$ RQM [Faustov:2023phi]
$(4.3\pm 0.9)$ $\times~10^{-10}$ NRQCD/NRCSM [Gao:2022iam]
$2.3$ $\times~10^{-9}$ LC+LCDA [Kartvelishvili:2008tz]
$\rm\Upsilon(2S)$ + $\rm\Upsilon(2S)$ $(1.0\pm 0.2)$ $\times~10^{-10}$ NRQCD/NRCSM [Gao:2022iam]
$\rm\Upsilon(3S)$ + $\rm\Upsilon(3S)$ $(5.7\pm 1.2)$ $\times~10^{-11}$ NRQCD/NRCSM [Gao:2022iam]
@@
Experimentally, the double-$(\mathrm{Q}\overline{\mathrm{Q}})$ and $(\mathrm{Q}\overline{\mathrm{Q}})(\mathrm{Q}\overline{\mathrm{Q}}^{\prime})$ decays have been searched for at the LHC [CMS:2022fsq], but the current $\mathcal{O}(10^{-3}$$10^{-4})$ limits are 4–5 orders-of-magnitude larger than the predictions. Their production at the HL-LHC, as well as at any future lepton collider, appears unfeasible, and only a machine like FCC-hh will have the Higgs production rates required to reach those decay modes, as indicated in Fig. 12.
@@ IV.1 Rare three-body Z boson decays
The two-body decay of the Z boson into two massless vector particles ($\rm Z\to\gamma\gamma,gg$) is forbidden by the Landau–Yang theorem, which states that a spin-1 boson cannot decay into two massless bosons [Landau:1948kw, Yang:1950rg], and of course the $\rm Z\to\gamma g$ decay is forbidden by conservation of colour. On the other hand, three-body decays into gauge bosons, or into a photon plus two neutrinos, are possible but can only be induced at the loop level in the SM (Fig. 13). Calculations for these virtual processes exist for many years [Laursen:1982rg, Glover:1993nv, Hopker:1993pb, Hernandez:1999xn] but they are not all always consistent among each other and, to our knowledge, have not been revised since then, although key underlying parameters, such as the $\alpha_{\text{s}}(m_{{}_{\mathrm{Z}}})$ coupling constant, hadronic Z boson partial widths, and top quark mass, have (noticeably) changed. For example, the original $\rm Z\to 3\gamma$ calculations from [Glover:1993nv] used $m_{\mathrm{t}}=124$ GeV, whereas the calculation of the $\rm Z\to 3g$ decay of [Hopker:1993pb] used the $m_{\mathrm{b}}\to 0$, $m_{\mathrm{t}}\to\infty$ limits, and the prediction for $\rm Z\to g+g+\gamma$ from [Laursen:1982rg] used $\alpha_{\text{s}}(m_{{}_{\mathrm{Z}}})=0.17$, $m_{\mathrm{t}}=20$ GeV, and $\Gamma_{\mathrm{Z}}=\Gamma_{\mathrm{Z\to\mathrm{q}\overline{\mathrm{q}}}}\approx 2$ GeV. We note also that many of these older works, including the review of Ref. [Perez:2003ad], provide contradictory values for some of these different branching fractions. We have therefore recalculated the rates for all these Z boson rare decays with MG5_aMC in the SM framework with QCD and EW loop corrections computed up to NLO accuracy, with massive fermions in the loops where needed, using the input parameters of Table 2. The updated branching fractions (Table 10) change with respect to the existing ones (using the old formulas with updated SM parameters, where needed) by 10–25% up or down, except for the $\rm Z\to\gamma+\nu\overline{\nu}$ case, where our result is six times smaller than that estimated in [Hernandez:1999xn]. This older work quoted a partial width for the triangle diagram which is smaller than our results for this contribution, whereas the relative impact of box diagrams was not clearly specified. So we suspect that a larger cancellation of contributions is present in our setup compared to that in this previous study.
@@
Branching fraction Framework 2024 HL-LHC FCC-$e$
[…]
$1.8$ $\times~10^{-6}$ loop SM, MG5_aMC (this work)
$\rm Z\;\to\;g+g+\gamma$ $7.0$ $\times~10^{-7}$ NLO SM [^9] With updated SM parameters. The original result was $\mathcal{B}=1.8\times 10^{-6}$ [Laursen:1982rg] [Laursen:1982rg]
$6.9$ $\times~10^{-7}$ loop SM, MG5_aMC (this work)
$\rm Z\;\to\;g+g+\gamma_{\mathrm{soft}}$ $1.2$ $\times~10^{-9}$ loop SM, MG5_aMC (this work)
[…]
$6.6$ $\times~10^{-10}$ loop SM, MG5_aMC (this work)
[…]
$\rm Z\;\to\;\gamma+\nu+\hat{\nu}$ $7.2$ $\times~10^{-10}$ loop SM [Hernandez:1999xn]
$1.4$ $\times~10^{-10}$ loop SM, MG5_aMC (this work)
@@
Figure 14: Branching fractions (in negative log scale) of rare three-body Z boson decays. The most recent theoretical predictions (blue bars) are compared to current experimental upper limits (violet) and expected conservative HL-LHC bounds (orange). The red vertical line indicates the expected FCC-$e$ reach based only on the total number of Z bosons to be produced.